obtain the differential equation associated with the given primitive, a and b : y = tan (x +a)
1
Expert's answer
2020-09-15T15:58:27-0400

The given equation is

y=tan(x+a)

we can write the above equation as

"tan^{-1}y=x+a"

differentiate both sides with respect to x

"\frac{1}{1+y^2}\frac{dy}{dx}=1"


"\frac{dy}{dx}=1+y^2"


"y'-y^2-1=0"